f(x) = (x⁴+3x³-2x²-x+7) / x-5, when x = 5?


1
Expert's answer
2021-12-07T06:40:19-0500

"f(x) = \frac{(x^4+3x^3-2x^2-x+7)}{x-5}"

At "x=5"

"f(5)=\frac{(5^4+3(5)^3-2(5)^2-5+7)}{5-5}=\infty"